OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN The position of an object at time t is given by s(t) = -9 - 5t. Find the instantaneous velocity at t = 4 by finding the derivative.

OpenStudy (mathmale):

Given that s(t)=-9-5t, find s '(t) (the derivative). Know how to differentiate -9? -5t? If so, share your results here.

OpenStudy (mathmale):

If not, what do you need help with?

OpenStudy (quantummechanics):

\[v(t) = \frac{ ds(t) }{ dt }\]

OpenStudy (quantummechanics):

So as @mathmale as requested, taking the first derivative will give you the velocity, you have done derivatives right?

OpenStudy (anonymous):

not really, this question is just a worksheet for the beginning of a chapter to see what know, but I am actually curious how to solve it

OpenStudy (mathmale):

Keep this in mind: 1) The derivative of a constant is zero. 2) The derivative of a power of x, such as x^n, is\[\frac{ nx ^{n-1} }{ }\]

OpenStudy (anonymous):

Sure we can try that way

OpenStudy (quantummechanics):

Mhm, well in that case you should read up on derivatives, have you heard of the derivative using a limit definition? It requires quite a bit of information to tackle this problem, but if you want to know how to do this with no context, we can apply the power rule, \[\frac{ d }{ dx } x^n = nx^{n-1} \] and then substitute t=4 once we have the velocity.

OpenStudy (mathmale):

Abbs....would you please try doing this yourself, using the rules given you?

OpenStudy (mathmale):

\[\frac{ d }{ dx }x ^{1}=\frac{ d }{ dx }x=\frac{ dx }{ dx }=?\]

OpenStudy (anonymous):

KK I am trying it now

OpenStudy (mathmale):

That's "the derivative of the first power of x."